# Ziwei Shi1,2     Xiaoran Zhang1,2     Yan Xia3,4     Yu Zang1,2     Siqi Shen1,2     Cheng Wang1,2

### v2

@@ Abstract
We tackle the challenge of LiDAR-based place recognition, which traditionally depends on costly and time-consuming prior 3D maps. To overcome this, we first construct XA-L&RSI dataset, which encompasses approximately $110,000$ remote sensing submaps and $13,000$ LiDAR point cloud submaps captured in urban scenes, and propose a novel method, L2RSI, for cross-view LiDAR place recognition using high-resolution Remote Sensing Imagery. This approach enables large-scale localization capabilities at a reduced cost by leveraging readily available overhead images as map proxies. L2RSI addresses the dual challenges of cross-view and cross-modal place recognition by learning feature alignment between point cloud submaps and remote sensing submaps in the semantic domain. Additionally, we introduce a novel probability propagation method based on particle estimation to refine position predictions, effectively leveraging temporal and spatial information. This approach enables large-scale retrieval and cross-scene generalization without fine-tuning. Extensive experiments on XA-L&RSI demonstrate that, within a $100km^{2}$ retrieval range, L2RSI accurately localizes $83.27\%$ of point cloud submaps within a $30m$ radius for top-$1$ retrieved location. We provide a video to more vividly display the place recognition results of L2RSI at https://shizw695.github.io/L2RSI/.
Figure 1: (Left) We propose L2RSI for cross-view LiDAR-based place recognition without 3D pre-built map. Given a point cloud submap representing the surroundings of the robot and the absolute orientation provided by a LiDAR and a magnetometer, L2RSI provides the most probable location in a large-scale city using high-resolution remote sensing imagery. (Right) Place recognition performance in different retrieval ranges. Notably, L2RSI achieved a recall ($\!<\!30m$) of over $80\%$ at Top-1 retrieval within a range of $100km^{2}$.
@@ 1 Introduction
Conditions Tasks
GPS Range $100km^{2}$
IMU Database >$60000$
Query LiDAR+MAG Distance <$30m$
Reference RSI $Recall@1$ >$80\%$
@@
  • The proposed method exhibits remarkable generalization ability, achieves $Recall@1$ ($\!<\!30m$) of $83.27\%$ within a retrieval range of $100km^{2}$, and demonstrates promising potential on cross scene generalization without fine-tuning.
@@ 3 Methodology
Fig. LABEL: shows the overview of the proposed L2RSI. We will introduce the Data preprocessing module in Sec. LABEL:. Sec. LABEL: givess the approach to train the global descriptor extractor for the query point cloud submaps and the remote sensing submaps. Sec. LABEL: introduces STPE, which aims to estimate the probability density of the current query position and refine these candidates.
@@ 4.1 Benchmark dataset
Figure 2: Dataset configuration. In XA-L&RSI 2(a), different colored rectangles highlight test sets from databases of varying sizes, all sharing the same query trajectory (green). Besides, the training trajectory (light blue) and discarded trajectory at the boundary (red) are annotated. In Oxford-L&RSI 2(b), the green trajectory is as the query for the test set, while the entire remote sensing imagery serves as the database.
XA-L&RSI. We purchase high-resolution remote sensing imagery covering Xiang’an District from AIRSAT Technology Group, with a resolution of $0.5m$. We extract submaps of $60m\!\times\!60m$ from the original remote sensing imagery at intervals of $40m$, resulting in a one-third overlap between adjacent submaps. And the submaps lacking road semantics in their $30m\!\times\!30m$ central region are filtered out.
XA-L&RSI Submap number Area ($km^{2}$)
PC RSI
Training Set 12194 47913 65
Test Set-S 862 3094 4
Test Set-M 862 6479 9
Test Set-L 862 12119 16
Test Set-G 862 63047 100
Oxford-L&RSI Submap number Area ($km^{2}$)
PC RSI
11-14-02-26 1697 2103 2.8
14-12-05-52 1671 2103 2.8
@@
[…]
@@
[…]
Oxford-L&RSI. Oxford Radar RobotCar is a famous urban scene dataset [1] for localization. The data is collected with a Velodyne HDL-32E LiDAR. The dataset includes multiple repeated traversals, each covering about $10km$, as shown in Fig. 2(b). Based on this, we supplement the database with the corresponding remote sensing imagery, covering an area of approximately $2.8km^{2}$. And we simulate the magnetometer by adding random noise within ±$\pm 15^{\circ}$ to the ground truth of orientation, as the official dataset does not include a magnetometer. The source and resolution of remote sensing imagery, as well as the data preprocessing process, are the same as those of the XA-L&RSI. The data of 11-14-02-26, 14-12-05-52 are used as the test set to evaluate the generalization ability of the proposed method without fine-tuning, see Table 2.
4.2 Evaluation criteria
4.3 Implementation details
The proposed L2RSI is implemented with Pytorch [7]. The network is trained with Adam optimizer [5]. All experiments are performed on Ubuntu $20.04.1$ with a single NVIDIA GTX $3090$ GPUs with 24GB of memory.
4.4 Results
4.4.1 Evaluation on test sets
@@
The proposed L2RSI is the first to perform cross-view LiDAR-based place recognition in large-scale urban scenes with high-resolution remote sensing imagery. Several state-of-the-art cross-view image retrieval networks are capable of handling preprocessed cross-modal semantic images, e.g. GeoDTR [15], Sample4Geo [3] and FRGeo [14]. Firstly, we benchmark them to evaluate the performance of our method in the test sets of XA-L&RSI with different size databases. The training settings on XA-L&RSI for all methods are consistent, except for the image size and descriptor dimensions, which follow the official configurations. The comparison is summarized in Table 3. L2RSI (w. STPE) significantly outperforms the baseline methods, reaches a $Recall@1$ of $88.93\%$ with a database of $4km^{2}$. As the retrieval range expands, L2RSI (w. STPE) demonstrates increased robustness, with the database area increased to $100km^{2}$ and only a $5.66\%$ degradation in $Recall@1$ and a $4.31\%$ degradation in $Recall@10$. In addition, we also introduced a comparison of two refinement strategies, including SuperGlobal reranking [8] and classical particle filtering (PF). The former focuses on refining individual queries and database features, leading to only marginal improvements in performance. For the latter, we apply its principles within the retrieval framework, initializing the number of particles to $120$. With the integration of sequence information, the performance of L2RSI (w. PF) shows considerable improvement. However, in challenging cross-view and cross-modal retrieval tasks, particles frequently decay, necessitating re-initialization and impeding the flow of sequence information. The proposed STPE substitutes filtering with aggregation, efficiently utilizing spatio-temporal information to substantially enhance the unstable performance of single query.
@@
Secondly, we directly test the model on the Oxford-L&RSI without any fine-tuning. Given the stark differences in equipment models, road widths, architectural styles, etc., this challenge marks the highest level of generalization ability. In the lower part of Table 3, the baseline methods fail completely, while L2RSI (STPE) achieves $Recall@1$ over $40\%$ and $Recall@10$ above $60\%$ on both trajectories. Although the current results are not yet sufficient for real-world deployment, they demonstrate great potential for future applications.

### v3

# L2RSI: Cross-view LiDAR-based Place Recognition for Large-scale Urban Scenes via Remote Sensing Imagery

@@ Abstract
We tackle the challenge of LiDAR-based place recognition, which traditionally depends on costly and time-consuming prior 3D maps. To overcome this, we first construct XA-L&RSI dataset, which encompasses approximately $110,000$ remote sensing submaps and $13,000$ LiDAR point cloud submaps captured in urban scenes, and propose a novel method, L2RSI, for cross-view LiDAR place recognition using high-resolution Remote Sensing Imagery. This approach enables large-scale localization capabilities at a reduced cost by leveraging readily available overhead images as map proxies. L2RSI addresses the dual challenges of cross-view and cross-modal place recognition by learning feature alignment between point cloud submaps and remote sensing submaps in the semantic domain. Additionally, we introduce a novel probability propagation method based on particle estimation to refine position predictions, effectively leveraging temporal and spatial information. This approach enables large-scale retrieval and cross-scene generalization without fine-tuning. Extensive experiments on XA-L&RSI demonstrate that, within a $100km^{2}$ retrieval range, L2RSI accurately localizes $83.27\%$ of point cloud submaps within a $30m$ radius for top-$1$ retrieved location. Our project page is publicly available at https://shizw695.github.io/L2RSI/.
Figure 1: (Left) We propose L2RSI for cross-view LiDAR-based place recognition without 3D pre-built map. Given a point cloud submap representing the surroundings of the robot and the absolute orientation provided by a LiDAR and a magnetometer, L2RSI provides the most probable location in a large-scale city using high-resolution remote sensing imagery. (Right) Place recognition performance in different retrieval ranges. Notably, L2RSI achieved a recall (<30m\!<\!30m) of over $80\%$ at Top-1 retrieval within a range of $100km^{2}$.
@@ 1 Introduction
Conditions
GPS
IMU
Query LiDAR+MAG
Reference RSI
Tasks
Range $100km^{2}$
Database >$60000$
Distance <$30m$
$Recall@1$ $80\%$
@@
  • The proposed method exhibits remarkable generalization ability, achieves $Recall@1$ (<30m\!<\!30m) of $83.27\%$ within a retrieval range of $100km^{2}$, and demonstrates promising potential on cross scene generalization without fine-tuning.
2 Related work
Uni-modal 3D place recognition. The early work by Uy et al. proposed PointNetVLAD [8], which utilized PointNet [9] as a point cloud feature extractor, followed by a NetVLAD layer [10] for global feature aggregation. LPD-Net [11] additionally introduced local information of points and aggregated global descriptors through GNN. SOE-Net [1] incorporated orientation encoding into the feature embedding process to capture spatial context and used a self-attention mechanism to aggregate global information. MinkLoc3D [12] applied sparse convolution and Generalized-Mean pooling [13] to extract discriminative features from sparse voxel representations of point clouds, and its extension, MinkLoc++ [14], further incorporated monocular camera images. SVT-Net [15] extended the transformer network to sparse voxels for capturing long-range contextual features. Zhou et al. [16] encoded geometric information by representing point clouds as normal distribution transformation cells. CASSPR [17] leveraged a dual-branch hierarchical cross-attention mechanism to integrate the advantages of point-based and voxel-based features. Crossloc3d [18] proposed a ground LiDAR place recognition method using aerial LiDAR as the database. However, such methods require an available and reliable prior 3D map, which is a demanding requirement in practice.
Cross-modal 3D place recognition. Cross-modal methods enhance usability by replacing costly prior 3D maps with data from alternative modalities. The work by Lee et al. [19] was the first to achieve LiDAR-based place recognition within a street-view image database. They created depth images from short sequence images using MantDepth [20] and established a consistent data representation with LiDAR range images through contrastive learning. Lip-Loc [21] aligned LiDAR point clouds with RGB camera images through range image projection. C2L-PR [22] achieved modality alignment by 2D to 3D mapping via depth estimation and semantic segmentation, and mitigates the FOV discrepancy through orientation voting. VXP [23] incorporated similarity constraints on local descriptors and established a voxel-pixel shared feature space through a two-stage training process. Text2Pos [24] and Text2Loc [25] determined the most probable location using textual descriptions. UniLoc [26] designed a universal place recognition working with any single query modality. Some researchers focused on publicly available maps, such as OpenStreetMap (OSM). Ruchti et al. [27] relied exclusively on road information extracted from LiDAR point clouds, associating it with the OSM road network to compute weights for Monte Carlo localization. Cho et al. [4] designed a hand-crafted rotation-invariant descriptor to retrieve the most similar location from the OSM database based on the relative position of LiDAR and buildings. The works most relevant to our research are a series of studies based on overhead images. Tang et al. [2] utilized ray-tracing to generate pseudo-point clouds from LiDAR scans and overhead images that are more similar to each other. In subsequent work [3], they introduced CycleGAN [28] to generate synthetic LiDAR images from satellite images and selected pseudo-pairs from sequential data to learn a shared embedding space. The place recognition range of the aforementioned methods is limited to known or predefined routes within urban scenes. In contrast, our L2RSI can be easily extended to unknown areas spanning tens to hundreds of square kilometers, offering broader application potential.
Sequence-Constrained Place Recognition. Garg et al. [29] and Malone et al. [30] introduced temporal information by training a sequence-matching network to enhance place recognition performance. However, remote sensing imagery is inherently static and lacks temporal continuity or directional cues, and such methods are not applicable. Prticle filtering [27, 31] is widely used in localization and tracking tasks, where it recursively refines the particles based on the motion model of the system, gradually converging them towards the position of the robot. The main difference from the proposed STPE is that our approach is particle aggregation rather than filtering. This is due to the inherent content discrepancies between remote sensing imagery and LiDAR data, where the observation information from a single query is often highly unreliable. Such toxic queries can lead to the filtering out of most, if not all, trustworthy particles. In contrast, the proposed method ensures the smooth transmission of spatio-temporal information and is more suitable for extremely challenging place recognition.
Cross-view 2D place recognition. Since our method unifies cross-modal data into semantic images, some image-based place recognition methods can also be applied. Early works such as CVUSA [32], CVACT [33], and VIGOR [34] established the main benchmarks in image place recognition between street view and satellite view. GeoDTR [35], used CNNs to extract global features and enforced the introduction of a Transformer to extract additional information through a counterfactual loss. Sample4Geo [36] introduced the off-the-shelf ConvNeXt-B as a feature extractor and designed a dynamic similarity sampling strategy to construct training batches. Recent FRGeo [37] proposes a feature reorganization module to effectively enhance feature aggregation for BEV images. In this work, we find that directly applying these methods can not work well in our task due to the domain gaps.
Figure 2: Overview of the proposed L2RSI. It consists of three modules: the data preprocessing module (left), the training framework for Global Descriptor Extraction network (middle) and the inference framework (right).
@@ 4 Methodology
Fig. 2 shows the overview of the proposed L2RSI. In the Data preprocessing module (Sec. 4.1), on one hand, we perform semantic segmentation on remote sensing imagery, which are then evenly divided into submaps to serve as the semantic database. On the other hand, we conduct semantic segmentation on point cloud submaps, which are subsequently compressed into BEV images. During training, we learn representations of BEV images and remote sensing submaps within a shared semantic space to overcome the modality gap (Sec. 4.2). During inference, we extract the global descriptor of the given query point cloud submap and retrieve a set of potential candidates from the semantic database. Then we apply STPE to estimate the probability density of the current query position and refine these candidates (Sec. 4.3).
4.1 Data preprocessing
Remote sensing imagery. Satellite and LiDAR data exhibit discrepancies in content and style due to differences in viewpoints, imaging principles, and acquisition times. In practice, humans typically assess whether the point cloud and remote sensing imagery represent the same location by identifying high-level semantic information. Inspired by this, we employ AIE-SEG developed by Alibaba [^1] https://engine-aiearth.aliyun.com/#/portal/analysis, which is a well-established universal segmentation model for remote sensing interpretation, to perform semantic segmentation on remote sensing imagery. Intuitively, we select roads, sidewalks, vegetation, and buildings as the semantics of interest. Subsequently, a square window with side length $d$ slides uniformly across the semantic image, extracting submaps to form the semantic database, as shown in the top-left corner of Fig. 2. Notably, considering that robots travel along the road, we filter out submaps that are located far from the road label.
Point cloud. For the acquired short sequences of LiDAR scans, we build the point cloud submap using FastGICP [38]. Then, we introduce SphereFormer [39] to perform segmentation based on the same semantics as remote sensing imagery. To adapt to data collected by different LiDAR systems, we remove the intensity feature and retrain the model on SemanticKITTI [40]. While various unsupervised domain adaptation methods can potentially yield better semantic segmentation results, our approach is simple yet effective.
To obtain the same perspective as the remote sensing imagery, we then compress the point cloud into a bird’s-eye view (BEV) image, covering an area of approximately $d\times d$. The topmost points are retained to preserve the features that are commonly visible in the remote sensing imagery. Based on the orientation measured by the magnetometer, we rotate the BEV image to align the north direction upwards before feeding it into the global descriptor extraction network.
4.2 Semantic contrastive learning
Recently, several works focused on building cross-modal representations have emerged prominently across various zero-shot computer vision tasks, such as classification [41], recognition [42, 43, 44] and retrieval [45, 46]. Inspired by their work, we design a network that aligns cross-modal semantic features in a unified feature space through contrastive learning.
Network architecture. As shown the middle of Fig. 2, we employ a dual-branch network to extract global descriptor for pairs of semantic images $\{Q_{i}\}_{i=1}^{N}$ from point clouds and $\{P_{i}\}_{i=1}^{N}$ from remote sensing imagery, where $N$ is the batch size. The distance between the centers of the paired images is constrained to be less than $d/2$. Each branch consists of a semantic encoder, a generalized mean (GeM) pooling, and a fully connected (FCN) layer in a sequential configuration. The semantic encoder supports initialization with various pre-trained foundation models. We use MAE [47] here. Then, GeM pooling and the FCN layer generate global descriptors. Since both branches operate within the semantic domain, we fully share their weights. In this work, we set $d=60m$.
Loss function. During training, we ensure that image pairs within a batch serve as negatives for each other and use the symmetric InfoNCE loss as the contrastive learning objective [36], formulated as follows:
$\mathcal{L}\!=\!-\!log(\!\frac{\!exp(\!f_{i}^{Q}\!\cdot\!f_{i}^{P}\!/\!\tau\!)}{\!\sum_{j\in N}\!exp(\!f_{i}^{Q}\!\cdot\!f_{j}^{P}\!/\!\tau\!)}\!)\!-\!log(\!\frac{\!exp(\!f_{i}^{P}\!\cdot\!f_{i}^{Q}\!/\!\tau\!)}{\!\sum_{j\in N}\!exp(\!f_{i}^{P}\!\cdot\!f_{j}^{Q}\!/\!\tau\!)}\!),$ (2)
where $f_{i}^{Q}$ and $f_{i}^{P}$ represent the global descriptors of $Q_{i}$ and $P_{i}$, respectively. The temperature coefficient $\tau$ is set to $0.1$. Compared to the traditional triplet loss [8, 2, 3], the advantage is that increasing the number of negatives in a training batch greatly accelerates the training process.
4.3 Spatial-Temproal Particle Estimation
During inference, we first extract the global descriptor of the query point cloud submap $Q_{t}$ at time step $t$. We then calculate its similarity distance to the image submaps in the remote sensing database for determining the Top-C retrieval results. ( see the top-right corner of Fig. 2). Note that the global descriptors are extracted offline to improve retrieval efficiency.
To alleviate the unreliability of individual query retrieval results, we consider the query sequence $\{Q_{j}\}_{t-L+1}^{t}$ from time step tL+1t\!-\!L+\!1 to $t$. The retrieval results of $Q_{j}$ are treated as particles, where the state of each particle, denoted as $\textbf{x}_{j}^{[k]}$ with $k\in{1,2,...,K}$, refers to the 2D coordinates of candidate submaps in the remote sensing database. The state of the $k$-th particle of $Q_{j}$ is transitioned from time step t1t-1 to $t$ through the following function:
$\textbf{x}_{t}^{[k,j]}=M_{t-1}(\textbf{x}_{t-1}^{[k,j]})+\textbf{v}_{t}^{[k,j]},$ (3)
where $M_{t-1}$ is the motion model from time step t1t-1 to $t$ and $\textbf{v}_{t}^{[k,j]}$ a noise term. Unlike traditional particle filtering, we propagate particles from multiple past time steps to current time step $t$ using the motion model, and then construct a probability density function at time step $t$ based on the combined particle distribution.
In practice, we obtain the relative positions between LiDAR scans from FastGICP [38], and correct the direction in the world coordinate system every $20$ meters using the magnetometer, which subsequently enables the determination of particle states at each time step. We observe that these particles at each time step tend to cluster around some locations with similar features, which can be formulated as multiple Gaussian distributions:
$P(x,y)\!=\!\sum_{m=1}^{M}\!A_{m}\!\cdot\exp\!\left(\!-\frac{(x\!-\!\mu_{xm})\!^{2}}{2\sigma_{xm}^{2}}-\frac{(y\!-\!\mu_{ym})\!^{2}}{2\sigma_{ym}^{2}}\!\right)\!,\!$ (4)
where $M$ represents the number of Gaussian distributions and we use DBSCAN clustering top-K candidates ($K<C$) to determine it with a radius of $r=30m$. Let $x$ and $y$ denote the coordinates of the particles. Since they are assumed to be independent, their covariance is 0. $A_{m}$ indicates the contribution of the $m$-th Gaussian distribution to the overall distribution, which is computed as follows:
$A_{m}=\frac{N_{m}}{\sum_{i=1}^{M}N_{i}},$ (5)
where $N_{i}$ is the number of candidates in the $i$-th cluster. The candidates in $m$-th cluster are represented as $\{(x_{i}^{m},y_{i}^{m})\}_{i}^{N_{m}}$. The center of Gaussian distribution $\mu_{xm}$ and $\mu_{ym}$ is the geometric center of $m$-th cluster:
$(\mu_{xm},\mu_{ym})=(\frac{1}{N_{m}}\sum_{i=1}^{N_{m}}x_{i}^{m},\frac{1}{N_{m}}\sum_{i=1}^{N_{m}}y_{i}^{m}),$ (6)
The standard deviations $\sigma_{xm}$ and $\sigma_{ym}$ are estimated as follows:
$\!\sigma\!_{xm}\!=\!\sqrt{\!\frac{1}{N_{m}}\!\!\sum_{i=1}^{N_{m}}(\!x_{i}^{m}\!-\!\mu_{xm}\!)\!^{2}},\sigma\!_{ym}\!=\!\sqrt{\!\frac{1}{N_{m}}\!\!\sum_{i=1}^{N_{m}}(\!y_{i}^{m}\!-\!\mu_{ym}\!)\!^{2}},$ (7)
Therefore, we estimate the probability density function of the query $Q_{t}$ by propagating the retrieval result distribution of the query sequence $\{Q_{j}\}_{t-L+1}^{t}$ from time step tL+1t\!-\!L\!+\!1 to $t$. Assuming the relative position between $Q_{j}$ and $Q_{t}$ is $(\Delta x_{j},\Delta y_{j})$. The distribution of the retrieval results of $Q_{j}$ after propagation according to (4) is:
$\tilde{P}\!_{j}(x,y)\!=\!\sum_{m=1}^{M}A_{m}\!\cdot\exp\!\left(\!-\frac{(x\!-\!\mu_{xm}^{j})\!^{2}}{2\sigma_{xm}^{2}}-\frac{(y\!-\!\mu_{ym}^{j})\!^{2}}{2\sigma_{ym}^{2}}\!\right)\!,\!$ (8)
where $\mu_{xm}^{j}\!=\!\mu_{xm}\!+\!\Delta x_{j}$ and $\mu_{ym}^{j}\!=\!\mu_{ym}\!+\!\Delta y_{j}$. This implies that the particle states at the current time step are derived from their estimated values at the previous step, rather than from the positions associated with the retrieved remote sensing images. The probability density function of the position at time step $t$ can be formulated as a Gaussian Mixture Model:
$P_{t}(x,y)=\frac{1}{L}\sum_{j=t-L+1}^{t}\tilde{P}_{j}(x,y),$ (9)
where the parameters of $\tilde{P}_{j}(x,y)$ is updated as the vehicle moves. Then, we assign a probability value to each remote sensing submaps in the database:
$P_{t}(u,v)=\frac{1}{S}\int_{u-r}^{u+r}\int_{v-r}^{v+r}P_{t}(x,y)\,dy\,dx,$ (10)
where $(u,v)$ is the center coordinate of the remote sensing submap. We take the average value within a rectangle with an area of $S\!=\!4r^{2}$ as the probability value for this remote sensing submap. To avoid cumulative errors, we only consider the query sequence of $250m$, and re-rank the retrieval results of $Q_{t}$ based on probability values.
@@ 5.1 Benchmark dataset
XA-L&RSI. We purchase high-resolution remote sensing imagery covering Xiang’an District from AIRSAT Technology Group, with a resolution of $0.5m$. We extract submaps of $60m\!\times\!60m$ from the original remote sensing imagery at intervals of $20m$, resulting in a two-third overlap between adjacent submaps. And the submaps lacking road semantics in their $30m\!\times\!30m$ central region are filtered out.
Oxford-L&RSI. Oxford Radar RobotCar is a famous urban scene dataset [5] for localization.
@@
Figure 3: Dataset configuration. In XA-L&RSI 3(a), different colored rectangles highlight test sets from databases of varying sizes, all sharing the same query trajectory (green). Besides, the training trajectory (light blue) and discarded trajectory at the boundary (red) are annotated. In Oxford-L&RSI 3(b), the green trajectory is as the query for the test set, while the entire remote sensing imagery serves as the database.
The data is collected with a Velodyne HDL-32E LiDAR. The dataset includes multiple repeated traversals, each covering about $10km$, as shown in Fig. 3(b). Based on this, we supplement the database with the corresponding remote sensing imagery, covering an area of approximately $2.8km^{2}$. And we simulate the magnetometer by adding random noise within $\pm 15^{\circ}$ to the ground truth of orientation, as the official dataset does not include a magnetometer. The source and resolution of remote sensing imagery, as well as the data preprocessing process, are the same as those of the XA-L&RSI. The data of 11-14-02-26, 14-12-05-52 are used as the test set to evaluate the generalization ability of the proposed method without fine-tuning, see Table 2.
@@
Submap Num. XA-L&RSI Oxford-L&RSI
Training Set Test Set-S Test Set-M Test Set-L Test Set-G 11-14-02-26 14-12-05-52
PC 12194 862 862 862 862 1697 1671
RSI 47913 3094 6479 12119 63047 2103 2103
Area ($km^{2}$) 65 4 9 16 100 2.8 2.8
5.2 Evaluation criteria
5.3 Results
5.3.1 Evaluation on test sets
Table 3: Performance comparison on XA-L&RSI and Oxford-L&RSI. $Recall@1$ and $Recall@10$ ($\%$) are reported. The best result is rendered in bold.
Methods $Recall@1$ / $Recall@10$ ($\%$) $\uparrow$
@@ 5.3.1 Evaluation on test sets
The proposed L2RSI is the first to perform cross-view LiDAR-based place recognition in large-scale urban scenes with high-resolution remote sensing imagery. Several state-of-the-art place recognition networks are capable of handling preprocessed cross-modal semantic images, such as cross-modal 3D place recognition (LIP-Loc [21]) and cross-view 2D place recognition (GeoDTR [35], Sample4Geo [36], and FRGeo [37]). Firstly, we benchmark them to evaluate the performance of our method in the test sets of XA-L&RSI with different size databases. The training settings on XA-L&RSI for all methods are consistent, except for the image size and descriptor dimensions, which follow the official configurations. Notably, we replace the LiDAR range image projection used in the original LIP-Loc with a BEV projection.The comparison is summarized in Table 3. L2RSI (w. STPE) significantly outperforms the baseline methods, reaches a $Recall@1$ of $88.93\%$ with a database of $4km^{2}$. As the retrieval range expands, L2RSI (w. STPE) demonstrates increased robustness, with the database area increased to $100km^{2}$ and only a $5.66\%$ degradation in $Recall@1$ and a $4.31\%$ degradation in $Recall@10$. In addition, we also introduced a comparison of two refinement strategies, including SuperGlobal reranking [48] and classical particle filtering (PF). The former focuses on refining individual queries and database features, leading to only marginal improvements in performance. For the latter, we apply its principles within the retrieval framework, initializing the number of particles to $120$. With the integration of sequence information, the performance of L2RSI (w. PF) shows considerable improvement. However, in challenging cross-view and cross-modal retrieval tasks, particles frequently decay, necessitating re-initialization and impeding the flow of sequence information. The proposed STPE substitutes filtering with aggregation, efficiently utilizing spatio-temporal information to substantially enhance the unstable performance of single query.
Table 4: Ablation study about effectiveness of components. The results include $Recall@1$ and $Recall@30$ ($\%$) on Test Set-S and 11-14-02-26. The best result is rendered in bold. Sem.: Semantic segmentation. STPE: Spatial-Temproal Particle Estimation. Orient.: Orientation.
Sem. STPE Orient. $Recall@1$ / $Recall@30$ ($\%$) $\uparrow$
Test Set-S 11-14-02-26
$\surd$ $\surd$ 50.43/67.90 8.13/25.42
$\surd$ $\surd$ 30.05/85.50 10.19/45.90
$\surd$ $\surd$ 54.37/76.14 4.92/21.24
$\surd$ $\surd$ $\surd$ 88.93/95.94 42.29/70.15
@@
Secondly, we directly test the model on the Oxford-L&RSI without any fine-tuning. Given the stark differences in equipment models, road widths, architectural styles, etc., this challenge marks the highest level of generalization ability. In the lower part of Table 3, the baseline methods fail completely, while L2RSI (STPE) achieves $Recall@1$ over $40\%$ and $Recall@10$ above $60\%$ on both trajectories. Although the current results are not yet sufficient for real-world deployment, they demonstrate great potential for future applications. Please refer to Sec. A.3 for comparative experiments in more scenes.
5.3.2 Ablation study
The following ablation studies evaluate the impact of different components and settings on the performance of L2RSI.
Semantic information. In the first row of Table 4, we bypass semantic images, directly extracting submaps from raw remote sensing imagery and projecting 3D coordinates to generate point cloud BEV images. Compared to the full L2RSI (last row), $Recall@1$ on test set-S decreases by $38.50\%$, and on Oxford-L&RSI, $Recall@1$ dropped to $8.13\%$. This demonstrates that semantics serve as reliable information capable of bridging domain differences, playing a pivotal role in enhancing the generalization ability.
Table 5: Ablation study about the sampling rate $\lambda$. $Recall@1$ ($\%$) on Test Set-S is reported. The runtime (ms) here only involves STPE. The best result is rendered in bold.
Test set-S $10\%$ $20\%$ $30\%$ $40\%$ $50\%$ $60\%$ $70\%$ $80\%$ $90\%$ $100\%$
$Recall@1$ ($\%$) $\uparrow$ 85.98 88.07 88.93 88.44 88.44 88.07 88.07 87.82 88.07 88.44
$RT(ms)$ $\downarrow$ 24.1 29.4 31.7 34.7 38.0 40.2 44.3 47.5 52.2 55.4
Figure 4: Ablation study about the number of queries in the sequence ($L$) and the number of particles for each query ($K$). Solid dots represent the optimal performance for each sequence length.
(a) retrieval results on Test Set-S
(b) retrieval results on 11-14-02-26
Figure 5: Visualization of retrieval results on XA-L&RSI 5(a) and Oxford-L&RSI 5(b). Points closer to red indicate that more retrieval results are required to achieve correct retrieval at that location. (I) shows the results without STPE, and (II) shows the results with STPE.
STPE. We ablate STPE in the second row of Table 4, and observe a performance drop of $58.88\%$ and $32.10\%$ in $Recall@1$ for the two test sets, respectively. In STPE, the GMMs are computed based on the top-K results (K=30K\!=\!30) retrieved from the query sequence. Therefore, we also provide $Recall@30$ as a reference.
To further explore the impact of key parameters in STPE, we experiment with different settings for the number of queries $L$ in the sequence and the number of particles $K$ for each query. The results are presented in Fig. 4. We observe that $L$ is positively correlated with retrieval performance, while $K$ exhibits a peak. Additionally, larger values of $L$ and $K$ generally lead to an increased runtime. Considering the trade-off between performance and computational cost, $L=50$ and $K=30$ are adopted as default configurations.
In order to reduce the runtime, we introduce the sampling rate $\lambda$ to decrease the number of queries involved in STPE, without changing the sequence length. As shown in Table 5, when $\lambda=30\%$, $Recall@1$ reaches $88.93\%$, while the refinement overhead per query is only $31.7ms$, demonstrating an excellent balance between performance and efficiency. This phenomenon is attributed to the fact that the adjacent point cloud submaps are spaced only $5m$ apart, resulting in highly redundancy. The sampled queries still encompass the majority of critical information. The qualitative results in Fig. 5 clearly illustrate the effectiveness of STPE.
5.3.3 Impact of Motion Model Failure
Table 6: Impact of motion model failure in STPE stage. Recall@1Recall@1($\%$) on Test Set-G are reported. The parentheses indicate the performance drop compared to the additional noise-free setting.
Recall@1Recall@1($\%$) $\epsilon_{yaw}=\pm 0^{\circ}$ $\epsilon_{yaw}=\pm 5^{\circ}$ $\epsilon_{yaw}=\pm 10^{\circ}$ $\epsilon_{yaw}=\pm 15^{\circ}$ $\epsilon_{yaw}=\pm 30^{\circ}$
$\epsilon_{xy}=\pm 0$m 83.27(-0.00) 80.57(-2.70) 78.72(-4.55) 60.76(-22.51) 48.46(-34.81)
$\epsilon_{xy}=\pm 1$m 82.77(-0.50) 80.20(-3.07) 77.85(-5.42)| 59.77(-23.50) 44.26(-39.01)
$\epsilon_{xy}=\pm 4$m 71.58(-11.69) 67.04(-16.23) 66.91(-16.36) 56.70(-26.57) 41.39(-41.88)
$\epsilon_{xy}=\pm 9$m 67.04(-16.23) 64.94(-18.33) 61.50(-21.77) 54.37(-28.90) 39.12(-44.51)
$\epsilon_{xy}=\pm 16$m 57.31(-25.96) 53.87(-29.40) 53.75(-29.52) 50.92(-32.35) 35.91(-47.36)
In STPE stage, dynamic objects such as pedestrians and vehicles may introduce disturbances into the motion estimation. To investigate the impact of progressive motion model degradation on place recognition performance, we introduce varying levels of yaw angle noise $\epsilon_{yaw}(^{\circ})$ and translational noise $\epsilon_{xy}(m)$ into the estimated motion model. The translation noise is defined by the noise in the x-direction and y-direction as follows:
$\epsilon_{xy}=\sqrt{(}\epsilon_{x}^{2}+\epsilon_{y}^{2}),$ (11)
where we take $\epsilon_{x}=\epsilon_{y}$ for convenience. The average Recall@1Recall@1($\%$) of three repeated experiments on Test Set-G ($100km^{2}$) of XA-L&RSI are shown in Table 6. Please note that the disturbances in the table represent intentionally introduced additional noise, based on the noise inherently caused by the FastGICP [38] algorithm and the magnetometer used in the system. $\epsilon_{yaw}=\pm 5^{\circ}$ and $\epsilon_{xy}=\pm 1$m indicates that, in addition to the inherent noise, random yaw angle noise within a range of $\pm 5^{\circ}$ and random translation noise within a range of $\pm 1m$ are introduced.
The results demonstrate that our method can tolerate a certain level of additional noise in the estimated motion model ($\epsilon_{yaw}<\pm 10^{\circ}$ and $\epsilon_{xy}<\pm 1$m), which is typically introduced by dynamic objects affecting the FastGICP algorithm. In practice, we find that excluding the points of the data collection vehicle itself from the raw point cloud effectively reduces this noise to an acceptable level.
5.3.4 Computational cost analysis
Table 7: Computational cost analysis on Test Set-S. The number of parameters (M) and runtime (ms) are reported. The best result is rendered in bold.
Metric GDE STPE
P (M) $\downarrow$ $RT(ms)$ $\downarrow$ P (M) $\downarrow$ $RT(ms)$ $\downarrow$
LIP-Loc [21] 46.09 5.7 / /
GeoDTR [35] 26.82 6.5 / /
Sample4Geo [36] 88.59 7.6 / /
FRGeo [37] 27.82 4.6 / /
L2RSI 86.58 5.8 / 31.7
In this section, we analyze the number of parameters and runtime of L2RSI in the global descriptor extraction and STPE stages. As reported in Table 7, since Sample4Geo and the proposed L2RSI use heavier backbones for feature extraction, i.e., convnext-B and MAE-B, they have more parameters. However, the runtime difference during the global descriptor extraction (GDE) is minimal. STPE significantly refines retrieval results, with an additional cost of $31.7ms$. Nevertheless, the total time of $37.5ms$ still allows for real-time operation (<100ms<100ms).